find the number such that when 14 is added to two thirds of it, the result is one and a quater times the original number.
Answer by stanbon(75887)

find the number such that when 14 is added to two thirds of it, the result is one and a quater times the original number.
------------------------
EQUATION:
(2/3)x + 14 = (5/4)x
Multiply thru by 12 to get:
8x + 12*14 = 15x
7x = 12*14
x = 12*2
x = 24
======================
Cheers,
Stan H.